How large a force is required to accelerate a 1600-kg car from rest to 25/ms over a distance of 200m?
Answer #1

m- 1600 kg Reitial velocity (o)= amis final velocity (f) = 25m/s distance (5) - 200m lale know that we = mz+as. 252 - 2ax 200